### v1

# ] 1NLPR, MAIS, CASIA  2UCAS  3ByteDance

Abstract
Models like OpenAI-o3 pioneer visual grounded reasoning by dynamically referencing visual regions, just like human “thinking with images”. However, no benchmark exists to evaluate these capabilities holistically. To bridge this gap, we propose TreeBench (Traceable Evidence Evaluation Benchmark), a diagnostic benchmark built on three principles: (1) focused visual perception of subtle targets in complex scenes, (2) traceable evidence via bounding box evaluation, and (3) second-order reasoning to test object interactions and spatial hierarchies beyond simple object localization. Prioritizing images with dense objects, we initially sample 1K high-quality images from SA-1B, and incorporate eight LMM experts to manually annotate questions, candidate options, and answers for each image. After three stages of quality control, TreeBench consists of 405 challenging visual question-answering pairs, even the most advanced models struggle with this benchmark, where none of them reach 60% accuracy, e.g., OpenAI-o3 scores only 54.87. Furthermore, we introduce TreeVGR (Traceable Evidence Enhanced Visual Grounded Reasoning), a training paradigm to supervise localization and reasoning jointly with reinforcement learning, enabling accurate localizations and explainable reasoning pathways. Initialized from Qwen2.5-VL-7B, it improves V* Bench (+16.8), MME-RealWorld (+12.6), and TreeBench (+13.4), proving traceability is key to advancing vision-grounded reasoning. The code is available at https://github.com/Haochen-Wang409/TreeVGR.
[…]
1 Introduction
Figure 1: Qualitative examples from TreeBench for each discipline. Each question requires focused visual parsing on mere objects, and some even request second-order reasoning beyond precise localization. Moreover, the bounding boxes of all target instances are provided, ensuring a traceable evaluation. All these questions are challenging, as OpenAI-o3 [43] and Gemini-2.5-Pro [10] cannot answer them correctly simultaneously.
Recent breakthroughs in Large Language Models (LLMs) reasoning, such as OpenAI-o1 [42] and DeepSeek-R1 [15] with remarkable test-time scaling properties, have motivated researchers to explore reasoning for Large Multimodal Models (LMMs) [18, 64, 65, 6]. These models are typically remarkable in their mathematical and scientific reasoning, particularly through text-space reasoning. However, they exhibit critical limitations when applied to perception-heavy tasks [19] or general multimodal benchmarks [62], primarily due to accumulated language bias from their exclusive reliance on textual reasoning pathways. A paradigm shift toward visual grounded reasoning emerged with models like OpenAI-o3 [43], which is able to “think with images” by dynamically referencing and amplifying task-relevant regions during reasoning, resulting in image-text interleaved reasoning pathways. Yet, despite growing interest, the community currently lacks comprehensive evaluation benchmarks for assessing these capabilities.
Classical benchmarks like POPE [26], MMBench [30], SEED-Bench [24], and MMMU [71] usually overlook fine-grained localization and verifiable reasoning chains. Others [66, 72, 63, 11, 56, 55, 73] partially address localization but lack traceability or complex reasoning: V* Bench [66] is restricted to simple spatial queries (e.g., “Is A left of B?”) and risks data contamination with COCO-derived images [27]; MME-RealWorld [72], HR-Bench [63], and document benchmarks [4, 39, 31] support high-resolution inputs but lack traceable evidence and second-order reasoning such as perspective shifts. In short, these benchmarks fail to adequately evaluate three key elements central to visual grounded reasoning: nuanced visual grounding, traceable multi-step reasoning, and dynamic cross-modal interaction through interleaved box-text reasoning pathways.
To bridge this gap, we propose TreeBench (Traceable Evidence Evaluation Benchmark), designed around three foundational principles essential for evaluating true “thinking with images” capabilities:
  • Focused Visual Perception. It evaluates a model’s ability to identify subtle targets within cluttered, real-world scenes using detailed, precise, and unique textual descriptions, which requires hierarchical scene understanding and the discrimination of extremely similar distractors.
  • Traceable Evidence. It not only evaluates the final accuracy but also pioneers quantifiable evaluation of reasoning chains, resulting in an explainable, reliable, and transparent evaluation.
  • Vision-Centric Second-Order Reasoning Capabilities. It moves beyond simple object localization and primitive “what/where” queries. It focuses on complex physical interactions between objects (such as contact and occlusion), as well as spatial containment (inside/outside, above/below) and relative relationships with perspective transformation.
To conduct TreeBench, we sample 1K images from SA-1B [20], prioritizing images with dense objects, as SA-1B [20] offers high-resolution, real-world scenes with a large number of small and varied objects, making it particularly suitable for evaluating visual grounded reasoning. Subsequently, 8 experts with solid technical backgrounds are involved in hand-crafted annotation for 10 sub-tasks, as demonstrated in Figure 1. In particular, we present a semi-automated pipeline. Each of OpenAI-o3 [43] and Gemini-2.5-Pro [10] is required to create three distinct questions belonging to a specific subtask, accompanied by multiple-choice options and the respective correct answers. Subsequently, experts curated or replaced these to ensure quality and difficulty. We additionally incorporate a cross-verification stage for further quality control. Finally, TreeBench incorporates 405 high-quality and extremely challenging VQA pairs with accurate bounding boxes of target instances. A comprehensive comparison between our TreeBench and other related benchmarks is provided in Table 1. Key advantages are summarized as follows:
  • Annotation Quality. Unlike benchmarks relying on LMM-generated labels such as MMT-Bench [69] and SEED-Bench [24], our expert-driven process ensures correctness and extreme difficulty. However, relying on models would inevitably introduce significant noise, compromising the quality of the annotations. On the contrary, our TreeBench is manually designed by 8 LMM experts, ensuring the annotation correctness and ensuring the difficulty of all questions.
  • Small Target Objects. All questions in TreeBench focus on extremely small objects in complex real-world scenes, where target instances occupy an average of 3.05% of the image.
  • Traceable Evidence Evaluation. Our TreeBench provides bounding box annotations of each target instance. It not only evaluates the final answer, but also reveals the quality of intermediate reasoning steps. Those predicted bounding boxes serve as a window into its process, helping to diagnose the source of errors, i.e., whether the model misunderstood the question or failed to locate the relevant object.
  • Task Difficulty. While models approach saturation (>90%) on benchmarks like V* Bench [66], even open-sourced state-of-the-art performers like Qwen2.5-VL-72B [2] achieve only 42.2 on our TreeBench, implying a large potential improvement for future works.
Benchmark Resolution
Traceable Evidence
Annotation
Mean Area of
Target Objects ($\downarrow$)
Qwen2.5-VL-72B
Performance ($\downarrow$)
V* Bench [66] 2,246×1,583 85.9
HR-Bench-4K [63] 4,023×3,503 79.3
HR-Bench-8K [63] 5,727×4,430 76.0
MME-RealWorld [72] 2,076×1,434 62.9
TreeBench 2,152×1,615 3.05% 42.2
Table 1: Comparison between benchmarks related to “thinking with images”. TreeBench features traceable evidence annotations, as well as high input resolution and challenging questions. $\downarrow$ indicates the lower, the more challenging.
Figure 2: Normalized performance comparison with our TreeVGR and other works [2, 76, 50] on our TreeBench for each category.
Beyond evaluation, we further introduce TreeVGR (Traceable Evidence for Visual Grounded Reasoning), a training paradigm enhancing localization-driven visual reasoning. Previous attempts like [60, 76, 5, 13, 46, 44, 50, 32] solely supervise final answers and neglect intermediate region-of-interest generation processes. It becomes hard to quantify the actual contribution of the “grounding-then-answering” framework. On the contrary, we propose TreeVGR, a novel training methodology emphasizing traceable evidence through reinforcement learning (RL), which explicitly supervises bounding box generation.
Building on RL with conventional accuracy-based and formatting rewards, TreeVGR leverages a novel dual IoU reward to ensure both precision and recall in localizing the ground-truth bounding boxes for each target instance. To implement this, we curate 37K samples for RL training, each comprising an image, a question, an answer, and corresponding bounding box annotations for all target instances. Empirically, initialized from Qwen2.5-VL-7B [2], TreeVGR brings significant improvements on various benchmarks, i.e., +16.8 on V* Bench [66], +12.6 on MME-RealWorld-Lite [72], and +13.4 on our TreeBench. Moreover, as illustrated in Figure 2, compared with related approaches, our TreeVGR enables traceable and explainable reasoning pathways with more accurate localizations (mIoU), and finally contributes to bootstrapped overall performance.
In conclusion, TreeBench pioneers the evaluation of how models “think with images”, while TreeVGR establishes a blueprint for training them. Together, they significantly advance the depth and utility of multimodal reasoning assessment with traceable evidence.
2 Related Work
Large Multimodal Models. Initial breakthroughs in Large Multimodal Models (LMMs), such as Flamingo [1] and BLIP-2 [25], achieved this by integrating visual features into the LLM backbone via cross-attention mechanisms. A significant shift towards efficiency emerged with LLaVA [28], which introduced a much more efficient approach. It projects visual features from a pre-trained encoder (e.g., CLIP [45]) directly into the LLM’s semantic space using a simple two-layer MLP. This paradigm of feature projection catalyzed rapid advancement. Subsequent research has dramatically scaled LMM capabilities and tackled increasingly complex tasks [57, 29, 23, 58, 2, 61, 22, 77, 67, 59]. A critical frontier has been handling high-resolution inputs. Models like LLaVA-NeXT [29] and InternVL-1.5 [8] adopt any resolution strategy. Qwen2-VL [61] and Qwen2.5-VL [2] introduce multimodal Rotary Position Embedding (mROPE) to support arbitrary resolution inputs. Beyond resolution, scaling pretraining with high-quality data is also vital, as demonstrated by InternVL3 [77]. Collectively, these models represent the state-of-the-art, forming robust baselines for diverse real-world multimodal applications. Our work builds upon these advances by leveraging their strong native visual grounding capabilities. However, existing LMMs do not naturally perform an explicit "grounding-then-answering" process, often resulting in misaligned or incomplete responses. By explicitly modeling this sequential process, our approach ensures more accurate and interpretable answers through grounded reasoning.
Reasoning LMMs. The groundbreaking reasoning capabilities of LLMs, exemplified by systems like OpenAI-o1 [42] and DeepSeek-R1 [15] have motivated efforts to extend similar competencies to multimodal settings using reinforcement learning (RL) [51]. Early approaches primarily focused on equipping LMMs to solve complex math and science problems involving image inputs [18, 64, 65, 6]. Other approaches [48, 33, 3] directly adopt GRPO [47] to open-ended visual grounding. Moreover, some attempts [34, 40, 46, 44] focus on regions-of-interest localization before actually answering the question. A recent milestone, OpenAI-o3 [43], advanced multimodal reasoning by enabling dynamic image manipulation, e.g., cropping and zooming into regions of interest, to emulate human-like "thinking with images." Subsequent research has sought to replicate this capability through diverse strategies: constructing SFT data [60], vanilla RL [13], framing grounding as a function [76], decoupling grounding and answering [5], multi-task reinforcement learning [32], and curiosity-driven reasoning [50]. Critically, these RL-based methods supervise only the final answer. In contrast, our TreeVGR emphasizes traceable evidence during RL training, i.e., supervising generated bounding boxes to ensure precise localization throughout the reasoning process. By doing so, TreeVGR enables more transparent, reliable, and fine-grained control over the reasoning pipeline.
Benchmarks for LMMs. Current benchmarks lack comprehensive evaluation of multimodal models’ ability to “think with images”, a capability demanding three core competencies: (1) focused visual perception (identifying small targets in large scenes), (2) traceable evidence (evaluating generated bounding boxes for explainability), and (3) second-order reasoning (deriving insights beyond precise instance localization). Some benchmarks may partially satisfy the first condition. While some benchmarks address isolated aspects, critical gaps persist. Classical benchmarks like POPE [26], MMBench [30], SEED-Bench [24], and MMMU [71] usually overlook fine-grained localization and verifiable reasoning chains. V* [66] evaluates detailed attributes and spatial relationships (e.g., “Is A left of B?”) but relies on COCO-derived images [27], introducing high contamination risk. MME-RealWorld [72] and HR-Bench [63] support high-resolution inputs but lack traceable evidence, and their questions often become easy when grounded precisely. Crucially, no benchmark integrates all three requirements, particularly the need for complex reasoning conditional on precise grounding, e.g., perspective transform: “From the perspective of person A, what is the relative direction of object B?”. To bridge this gap, we propose TreeBench, the first benchmark designed explicitly for “thinking with images” with traceable, multistep evaluation. Beyond accuracy, TreeBench assesses: (1) region quality, i.e., faithfulness of generated regions-of-interest in visual reasoning chains, and (2) second-order reasoning, i.e., capabilities requiring inference beyond localization. State-of-the-art models, Gemini-2.5-Pro [10] and OpenAI-o3 [43], perform poorly on TreeBench (<60%), underscoring its rigor and the unmet challenges in multimodal reasoning.
3 TreeBench
TreeBench is designed to address a critical gap in multimodal evaluation by establishing the first comprehensive benchmark for assessing “thinking with images” capabilities. Specifically, it mainly evaluates (1) the ability of identifying small target objects with long, detailed, and unique text captions in large, complex, and real-world scenes, (2) the explainability of reasoning pathways and traceable evidence, and (3) second-order reasoning beyond precise localization. Our TreeBench systematically evaluates 10 core competencies through 405 distinct questions, organized into two progressive protocols, i.e., “Perception” and “Reasoning”, with representative examples demonstrated in Figure 1.
3.1 Task Definition
Perception evaluates the model’s ability to accurately “see” and “identify” specific content, which is one of the basic capabilities of directly extracting and interpreting visual information from every detail of the provided image. These tasks primarily evaluate first-order visual reasoning capabilities, where correct answers usually depend on the accurate localization of target questions (e.g., objects, regions, or text) and directly recognize their explicit attributes without requiring higher-level logical inference or abstract conceptualization. The success hinges on precise perceptual grounding. It includes:
  1. 1.
    Attributes evaluates the ability to identify and describe specific visual properties (e.g., color, shape, material, or precise classification) of objects or elements within images, particularly requiring attention to fine details, subtle distinctions, and accurate recognition of small-scale or context-dependent features.
  2. 2.
    Material measures the ability to analyze and distinguish material properties (e.g., texture, surface finish, composition, or physical state) through visual cues such as light reflection, transparency, wear patterns, or microscopic structural characteristics, requiring precise reasoning about tactile qualities and material-specific visual indicators.
  3. 3.
    Physical State assesses the ability to assess structural integrity (e.g., damage, wear, or breakage), detect positional states (e.g., open/closed, bent/straight), and interpret age-related features (e.g., freshness, decay) through precise analysis of visual cues like cracks, alignment anomalies, lighting/shadow patterns, or contextual degradation markers.
  4. 4.
    Object Retrieval probes the ability to interpret linguistically complex, spatially explicit descriptions and map them to visually subtle or contextually embedded targets in images, testing the integration of natural language understanding, spatial grounding, and discriminative object recognition under high specificity constraints.
  5. 5.
    OCR-Integrated Question-Answering evaluates the ability to extract text-based questions and answer options from images, requiring seamless integration of OCR, natural language understanding, and multimodal alignment to produce accurate responses grounded in both textual and visual modalities.
Reasoning evaluates the ability to analyze and infer meaningful conclusions beyond recognition. These tasks demand second-order visual reasoning capabilities, where correct answers require not only accurate localization but also higher-level cognitive operations over aggregated visual evidence. Precise perceptual grounding is just the first step for these tasks. It includes:
  1. 1.
    Perspective Transform measures the capacity to perform viewpoint transformations (e.g., aligning viewer-centric and agent-centric frames of reference) and interpret spatial relations under mirror-reversed or perspective-shifted conditions, testing the ability to disambiguate directional relationships that depend on the visualized entity’s orientation rather than the image’s literal pixel layout.
  2. 2.
    Ordering evaluates the ability to analyze linearly ordered arrangements of objects (e.g., left-to-right, front-to-back, or depth-based sequences) and resolve ordinal relationships by integrating spatial context with discriminative feature recognition, requiring precise localization within continuous layouts and contextual comparison of positional cues (e.g., adjacency, centrality, or extremity) to answer questions dependent on sequential alignment and relative placement.
  3. 3.
    Contact and Occlusion measures the ability to analyze physical interactions between multiple objects (e.g., direct contact, occlusion layers, or shadow-based overlaps) and resolve ambiguities in object identification by leveraging spatial dependencies, requiring precise parsing of contact cues (e.g., alignment, boundary fusion), occlusion boundaries (e.g., partial/full coverage, layer stacking), and contextual constraints to answer questions that hinge on understanding how objects physically coexist and obscure one another in complex scenes.
  4. 4.
    Spatial Containment benchmarks the ability to analyze hierarchical spatial relationships (e.g., containment, surface attachment, or regional boundaries) by parsing visual cues like object boundaries, spatial context, and contextual containment rules, requiring precise interpretation of containment hierarchies, surface dependencies, and regional constraints to resolve questions dependent on explicit spatial membership rather than isolated positional attributes.
  5. 5.
    Comparison assesses to compare attributes across multiple objects (e.g., distance, size, color) and resolve spatial or perceptual differences, requiring precise parsing of attribute discrimination and contextual distance estimation to answer questions demanding explicit comparison of visually co-present entities.
3.2 Annotation Pipeline
The TreeBench dataset was constructed through a systematic pipeline combining automated sampling, LMM-assisted generation, and three rounds of human validation. The annotation team contains eight human experts in LMMs, including six Ph.D candidates and two senior research scientists.
Image Selection. A total of 1K images are initially sampled from the SA-1B [20], with deliberate prioritization of images containing high-density objects (e.g., scenes with overlapping or clustered items), as SA-1B [20] offers high-resolution, real-world scenes with a large number of small and varied objects, making it particularly suitable for evaluating visual grounded reasoning. To ensure balanced representation across categories, 100 images are initially allocated per category.
First Round Quality Control. The annotation team manually evaluates the relevance and quality of each image for its assigned category. This step is critical for addressing category-specific requirements, e.g., the “Ordering” category necessitates images with visually similar or repetitive objects for practical reasoning tasks. Following this review, 647 images meet the criteria for inclusion in subsequent stages.
Automated Question Generation. Question-option-answer trios are then generated using two advanced LMMs, i.e., OpenAI-o3 [43] and Gemini-2.5-Pro [10], each tasked with producing three diverse, high-quality questions per image. Prompts are designed to emphasize task-specific complexity and visual-semantic alignment.
Second Round Quality Control. Human experts then manually review all six model-generated questions per image. For each image, annotators selected the most semantically coherent and task-relevant question from the pool of six, prioritizing: (1) alignment with the target subtask, (2) avoidance of trivial or ambiguous object referring, and (3) clarity and unambiguous answerability. If none of the six questions met these criteria, annotators manually constructed a new question tailored to the image. This step ensured that only high-quality, human-vetted questions advanced to the next stage.
Difficulty Filtering. Questions deemed insufficiently challenging are removed through model-based consensus screening. Specifically, any question answered correctly by all four state-of-the-art vision-language models (Qwen2.5-VL-72B [2], InternVL3-78B [77], GPT-4o [41], Gemini-2.5-Flash [9]) was excluded to ensure the benchmark retained meaningful difficulty.
Third Round Quality Control. The final cross-verification phase engages independent human annotators to cross-validate the accuracy and relevance of each question-option-answer pair. The final dataset comprised 405 rigorously validated questions.
Figure 3: Distribution of each discipline in TreeBench, which prioritizes reasoning over perception.
Figure 4: The ground-truth distribution of TreeBench with 3 instances of E and 4 instances of F.
Figure 5: Distribution of the number of instances in TreeBench, with one question with 8 target instances.
3.3 Statistics
Distribution of Each Subtask. As demonstrated in Figure 5, TreeBench emphasizes advanced reasoning tasks, accounting for 63% of the total subtasks (256 questions), while basic perception-related tasks constitute 37% (149 questions). Within the reasoning category, key subtasks reflect a focus on complex spatial and relational understanding. This structure underscores a deliberate prioritization of higher-order reasoning over foundational perceptual tasks, aligning with the goal of challenging models to process nuanced relationships and transformations rather than mere object recognition or attribute detection.
Distribution of Answers. As illustrated in Figure 5, the ground-truth distribution of TreeBench is dominated by four main categories: A (28.6%, 116 instances), B (27.9%, 113 instances), C (27.9%, 113 instances), and D (13.8%, 56 instances). These account for 98.2% of the total 405 instances. The remaining 1.8% (7 instances) includes E (3 instances) and F (4 instances). This structure highlights a balanced emphasis on categories A, B, and C, with D as a notable secondary group, while E and F represent minor but distinct components.
Distribution of the Number of Target Instances. Figure 5 shows the distribution of the number of target instances per question. The majority of questions in TreeBench require identifying 1 or 2 target instances, accounting for 41.5% (168 questions) and 44.9% (182 questions) of the total, respectively. Questions requiring 3, 4, 5, or 6 targets constitute smaller fractions: 4.2% (17 questions), 4.0% (16 questions), 4.0% (16 questions), and 1.2% (5 questions), respectively. Notably, a single question (highlighted in gray) demands 8 target instances, representing an extreme case. Overall, 86.4% of questions focus on 1–2 targets, suggesting a balance between simplicity and complexity in task design while incorporating rare multi-target scenarios for comprehensive evaluation.
Figure 6: The histogram of mean target instance areas per question with a low average of 0.0305 (indicating small target instances).
Distribution of Target Instance Area. We compute the relative area for each target instance using its bounding box, i.e., $\text{area}=\frac{1}{HW}(y_{2}-y_{1})(x_{2}-x_{1})$, where $H$ and $W$ are the input resolution. Figure 6 is the histogram of the mean area for each question. It illustrates that the majority of target instances in TreeBench are extremely small, with a sharp peak near 0.0 and a long tail extending to larger areas (up to 0.7). The mean area across all questions is 0.0305, confirming that targets are predominantly tiny. Most questions (highest frequency bin) involve target instances with areas clustered around 0.0 to 0.05, while only a small fraction require identifying larger objects. This distribution highlights the importance of addressing challenging scenarios where small-scale object detection and reasoning are crucial, potentially compromising model performance due to limited visual information.
Figure 7: Training pipeline of TreeVGR, including (a) a cold-start initialization stage and (b) a reinforcement learning with traceable evidence post-training stage.
4 TreeVGR
In this section, we introduce our TreeVGR. Specifically, we leverage the native grounding capabilities of pre-trained LMMs and unlock visual grounded reasoning capabilities, i.e., localizing regions-of-interest first and answering the question next, through a two-stage training pipeline shown in Figure 7, i.e., cold initialization introduced in Section 4.1 and reinforcement learning with traceable evidence elaborated in Section 4.2.
Notably, our TreeVGR does not require actually replaying cropped images as previous approaches [60, 76, 50] do, as text-space grounding is already effective. It leads to much more efficient training and inference procedures.
4.1 Cold-Start Initialization
While end-to-end reinforcement learning (RL) has demonstrated validity by [76] for visual grounded reasoning (VGR) tasks, its practical deployment remains hindered by extreme computational demands. Specifically, DeepEyes-7B [76] requests RL training on 47K samples across 32 episodes, a process requiring 32 H100 (80GB) GPUs operating continuously for 48 hours. Such resource intensity creates barriers to iterative experimentation and broader accessibility.
To address these limitations, we investigate a computationally efficient alternative. Initial attempts revealed significant training inefficiencies when applying direct RL to VGR: models required extensive iterations to autonomously identify task-relevant visual regions before generating answers. This bottleneck motivates our adoption of a cold initialization strategy as illustrated in Figure 7a. Specifically, we introduce a supervised fine-tuning (SFT) phase using a curated dataset comprising multimodal samples: each sample includes an image, a question, reasoning trajectories with corresponding bounding boxes, and a final answer. This structured initialization ensures VGR capabilities are established prior to RL.
Data Construction. We base our supervised fine-tuning (SFT) dataset on VGR-158K [60], which provides pseudo-chain-of-thought annotations paired with bounding boxes for visual reasoning tasks. However, to align with the grounding capabilities of our base model (Qwen2.5-VL series [2]), which outputs absolute coordinates rather than the normalized coordinates (ranging from 0 to 1) used by LLaVA-NeXT [29] in [60], we perform coordinate system conversion. Specifically, for each bounding box, we transform normalized coordinates $[r_{x_{1}},r_{y_{1}},r_{x_{2}},r_{y_{2}}]$ into absolute coordinates via $[x_{1},y_{1},x_{2},y_{2}]=[Wr_{x_{1}},Hr_{y_{1}},Wr_{x_{2}},Hr_{y_{2}}]$, where $H\times W$ is the resolution of the input image. Next, we filter samples to prioritize complex reasoning pathways, retaining only entries with multiple bounding boxes (i.e., more than one box per reasoning trajectory). This yields 35K samples, as multi-box interactions demand stronger spatial-temporal reasoning compared to single-box tasks. Subsequently, we construct a reflective subset of 4.7K samples among them by introducing controlled perturbations: for each sample, we (1) inject a synthetic error by inserting a randomly generated incorrect bounding box into the reasoning sequence, and (2) append the meta-cognitive prompt “Wait, this box seems to be wrong” immediately afterward, resulting in our TreeVGR-SFT-35K. This design explicitly trains the model to detect and correct erroneous visual grounding, which is a critical skill for robust real-world deployment.
Optimization. Initialized from Qwen2.5-VL-7B-Instruct [2], we train TreeVGR-7B-CI (“CI” here stands for Cold Initialization) with 8 GPUs using LLaMA-Factory [74], where the AdamW optimizer [36] with a learning rate of 5e-6 and a global batch size of 256 is utilized. The learning rate is decayed following a cosine schedule [35] with a warmup ratio of 0.1.
4.2 Reinforcement Learning with Traceable Evidence
We proceed to reinforcement learning (RL) to refine reasoning trajectories through traceable evidence supervision as demonstrated in Figure 7b. Specifically, the bounding boxes generated are evaluated using a box intersection-over-union (IoU) reward, a precise and interpretable metric that measures the alignment between predicted and ground-truth regions. This reward ensures explicit accountability to human-annotated visual evidence, guiding the policy toward spatially accurate and logically coherent reasoning pathways.
Reward Design. The total reward consists of three parts: an accuracy reward $R_{\text{acc}}\in\{0,1\}$, a formatting reward $R_{\text{format}}\in\{0,1\}$, and a dual Intersection-over-Union (IoU) reward $R_{\text{IoU}}\in[0,1]$:
$R=R_{\text{acc}}+R_{\text{format}}+R_{\text{IoU}},$ (1)
where the accuracy reward assesses whether the final answer is correct. We utilize exact-matching for multiple-choice questions, and leverage an online reward model, i.e., Qwen2.5-72B-Instruct [52], to judge whether the prediction is correct given the question and the ground-truth answer. The formatting reward ensures the reasoning process and the final answer must be enclosed between <think> and </think>, and <answer> and </answer>, respectively. The dual IoU reward measures the quality of predicted boxes against ground-truths. Specifically, for $N$ predicted bounding boxes $\{\hat{\bm{b}}_{i}\}_{i=1}^{N}$, where $\hat{\bm{b}}_{i}=[\hat{x}_{1}^{i},\hat{y}_{1}^{i},\hat{x}_{2}^{i},\hat{y}_{2}^{i}]$ and $M$ ground-truths $\{\bm{b}_{k}\}_{k=1}^{M}$, where $\bm{b}_{k}=[x_{1}^{k},y_{1}^{k},x_{2}^{k},y_{2}^{k}]$, the dual IoU is an average of a recall term and a precision term.
$R_{\text{IoU}}=\frac{1}{2}(R_{\text{IoU}}^{\text{R}}+R_{\text{IoU}}^{\text{P}}),$ (2)
where the first term $R_{\text{IoU}}^{\text{R}}$ indicates the recall and the second term $R_{\text{IoU}}^{\text{P}}$ means the precision. Specifically, the recall term ensures that each ground-truth bounding box $\bm{b}_{k}$ is matched with at least one prediction.
$R_{\text{IoU}}^{\text{R}}=\frac{1}{M}\sum_{k=1}^{M}\text{IoU}\left[\{\hat{\bm{b}}_{i}\}_{i=1}^{N},\bm{b}_{k}\right],$ (3)
where $\text{IoU}\left[\{\hat{\bm{b}}_{i}\}_{i=1}^{N},\bm{b}_{k}\right]=\max_{i}\text{IoU}(\hat{\bm{b}}_{i},\bm{b}_{k})$ indicates the maximum IoU between all predictions $\{\hat{\bm{b}}_{i}\}_{i=1}^{N}$ and each ground-truth $\bm{b}_{k}$. Maximizing this term ensures each ground-truth $\bm{b}_{k}$ is matched with at least one prediction. However, we empirically find that the policy model tends to enumerate all possible boxes to obtain a larger recall. Therefore, we introduce a dual term, i.e., $R_{\text{IoU}}^{\text{P}}$, to ensure the precision and discourage “empty” boxes that do not match with any ground-truths:
$R_{\text{IoU}}^{\text{P}}=\frac{1}{N}\sum_{i=1}^{N}\text{IoU}\left[\{\bm{b}_{k}\}_{k=1}^{M},\hat{\bm{b}}_{i}\right].$ (4)
Similarly, $\text{IoU}\left[\{\bm{b}_{k}\}_{k=1}^{M},\hat{\bm{b}}_{i}\right]=\max_{k}\text{IoU}(\bm{b}_{k},\hat{\bm{b}}_{i})$ indicates the maximum IoU between all ground-truths $\bm{b}_{k}$ and each prediction $\{\hat{\bm{b}}_{i}\}_{i=1}^{N}$. Maximizing this term encourages each prediction $\hat{\bm{b}}_{i}$ to be matched with at least one ground-truth. Therefore, simultaneous optimization of both recall and precision eliminates the need for exhaustive enumeration of bounding boxes, thereby contributing to more accurate reasoning pathways.
Attributes
Material
Phy. State
Obj. Retr.
OCR
Per. Trans.
Ordering
Con. & Oc.
Spa. Cont.
Comparison
Overall mIoU Perception Reasoning
Private Models
Gemini-2.5-Flash-0520 [9] 45.9 48.3 53.9 69.6 68.8 75.0 15.3 19.3 56.1 72.4 43.2
GPT-4o-1120 [41] 46.9 51.7 61.5 65.2 43.8 69.1 18.8 38.6 48.8 72.4 43.2
Gemini-2.5-Pro-0605 [10] 54.1 51.7 61.5 56.5 75.0 83.8 20.0 36.8 65.9 86.2 54.6
o3-0416 [43] 54.8 69.0 69.2 65.2 68.8 79.4 22.4 38.6 61.0 86.2 50.0
Open-source General Models
LLaVA-OneVision-7B [23] 37.3 55.2 53.8 56.5 50.0 32.4 21.2 22.8 41.5 72.4 36.4
LLaVA-OneVision-72B [23] 40.5 62.1 53.8 65.2 62.3 36.8 12.9 28.1 53.7 65.5 47.7
Qwen2.5-VL-7B [2] 37.0 55.2 53.8 56.5 62.5 27.9 20.0 35.1 39.0 44.8 43.2
Qwen2.5-VL-32B [2] 42.5 51.7 53.8 69.6 62.5 54.4 16.5 33.3 46.3 62.1 38.6
Qwen2.5-VL-72B [2] 42.2 65.5 69.2 56.5 56.3 48.5 11.8 33.3 51.2 72.4 38.6
InternVL3-8B [77] 38.8 51.7 69.2 56.5 56.3 33.7 21.2 24.6 39.0 72.4 43.2
InternVL3-38B [77] 42.0 51.7 61.5 52.2 68.8 51.5 12.9 33.3 56.1 65.5 38.6
InternVL3-78B [77] 46.4 62.1 61.5 52.2 68.8 52.9 16.5 33.3 61.0 86.2 45.5
Open-source Visual Grounded Reasoning Models
DeepEyes-7B [76] 37.5 30.0 62.1 53.8 65.2 68.8 51.5 11.8 24.6 36.6 51.7 47.7
Pixel-Reasoner-7B [50] 39.0 35.7 58.6 61.5 65.2 50.0 48.5 14.1 31.6 39.0 44.8 40.9
TreeVGR-7B 50.4 44.0 65.5 53.8 82.6 68.8 63.3 22.4 36.8 61.0 69.0 45.5
$\Delta$ v.s. Qwen2.5-VL-7B $\uparrow$ 13.4 $\uparrow$ 11.7 – 0.0 $\uparrow$ 26.1 $\uparrow$ 6.3 $\uparrow$ 35.4 $\uparrow$ 2.2 $\uparrow$ 1.7 $\uparrow$ 22.0 $\uparrow$ 24.2 $\uparrow$ 2.3
Table 2: Selected results of different models on TreeBench. Evaluations of open-source general models are implemented using VLMEvalKit [12], while evaluations of visual grounded reasoning models are conducted by us. Reasoning pathways of o3 [43] are unavailable, and thus traceable evaluations are not valid. Best performances for open-source models are highlighted in bold. Our TreeVGR-7B achieves comparable performance with InternVL3-78B [77].
Data Construction. As discussed above, TreeVGR incorporates a novel dual IoU reward, which means each sample should contain ground-truth bounding boxes during the RL phase. To this end, we filter hard samples from the original 191K training set of V* [66] using Qwen2.5-VL-7B-Instruct [2], resulting in 30K samples. Additionally, we incorporate the VisDrone dataset [78], which is originally designed for detection and tracking under UAV images, which offers extremely high-resolution, real-world scenes with a large number of small and varied objects and their corresponding bounding box annotations. We reformulate the training set and the validation set into 38K multiple-choice counting problems, and only retain samples with the ground-truth number ranging from 5 to 10, contributing to the final 7K samples. Finally, our TreeVGR-RL-37K consists of 30K open-ended question-answering samples from V* [66] and 7K multiple-choice problems from VisDrone [78].
Optimization. Initialized from TreeVGR-7B-CI, we train our final TreeVGR-7B with 8 GPUs, with another 8 GPUs serving the reward model, i.e., Qwen2.5-72B-Instruct [52], using vLLM [21]. We adopt Group Relative Policy Optimization (GRPO) [47], which has been proved to be effective and efficient for diverse tasks. We have also tried DAPO [70], but we find it unstable compared with GRPO. Therefore, we simply utilize the original GRPO [47]. We implement using EasyR1 [75], which is a clean fork of veRL [49]. We train our TreeVGR-7B with 5 epochs on TreeVGR-RL-37K, which is significantly less than DeepEyes-7B [76] (which is trained on 47K samples with 32 epochs).
5 Experiments
Figure 8: Distribution of IoU for each question in TreeBench.
Figure 9: Distribution of IoU for each question in TreeBench-Perception.
Figure 10: Distribution of IoU for each question in TreeBench-Reasoning.
Figure 11: Performance decoupling with AI2D [17].
Figure 12: Performance decoupling with MathVista [37].
Figure 13: Performance decoupling with MMStar [7].
5.1 Main Properties on TreeBench
Results on TreeBench. We evaluate four state-of-the-art private models, including GPT-4o-1120 [41] and o3-0416 [43] from OpenAI, and Gemini-2.5-Flash-0520 [9] and Gemini-2.5-Pro-0605 [10] from Google DeepMind. Additionally, 8 representative open-source general models are incorporated, including LLaVA-OneVision series [23], Qwen2.5-VL series [2], and InternVL3 series [77]. Furthermore, two very recent visual grounded reasoning models are also included, i.e., DeepEyes [76] and Pixel-Reasoner [50], as both of them follow a “grounding then answering” pipeline, with the capability of “thinking with images”. Table 2 presents per per-category performance of different models. Overall, OpenAI’s o3-0416 [43], the state-of-the-art visual grounded reasoning model, demonstrates the strongest perception abilities, as expected. Larger models usually contribute to better performance, while Qwen2.5-VL-32B [2] and its 72B counterpart as an exception. Notably, our TreeVGR-7B even achieves comparable performance with InternVL3-78B [77], demonstrating the effectiveness of the visual grounded reasoning pipeline.
Correlation between Localization and Performance. Importantly, for visual grounded reasoning models, our traceable evaluation demonstrates a positive correlation between localization precision and the overall performance, as illustrated in Table 2. This positive correlation between precise localization (mIoU) and overall performance is evident in the progressive improvement from DeepEyes-7B [76] to Pixel-Reasoner-7B [50] to our final TreeVGR-7B. As mIoU increases, the overall scores rise correspondingly, with TreeVGR-7B achieving the highest mIoU and strongest overall performance at the same time.
Beyond global analysis, we further plot the histogram of IoU for each question in Figure 10, where blue bars represent wrong predictions and orange bars are correct predictions. Overall, wrong predictions tend to have smaller IoU values. However, by going deeper through the lens of perception and reasoning, the relationship between mIoU and performance diverges. Precise localization (mIoU) aligns closely with perception performance demonstrated in Figure 10. In contrast, as shown in Figure 10, reasoning performance reveals a weaker correlation with mIoU, as improvements in localization alone fail to fully translate to complex reasoning tasks. This disconnect suggests that reasoning questions of TreeBench require second-order cognitive capabilities that go beyond precise spatial localization.
V* Bench [66] HR-Bench-4K [63] HR-Bench-8K [63]
Overall Attr. Spatial Overall Single Cross Overall Single Cross
Private Models
GPT-4o-1120 [41] 66.0
o3-0416 [43] 95.7
Open-source General Models
LLaVA-OneVision-7B [23] 70.7 73.0 60.5 64.3 74.8 53.8 59.8 65.3 54.3
LLaVA-OneVision-72B [23] 73.8 80.9 63.2 66.3 76.5 56.0 60.9 68.8 53.0
InternVL3-8B [77] 72.3 73.0 71.1 70.8 79.3 62.3 62.0 64.3 59.8
InternVL3-38B [77] 77.5 77.4 77.6 76.3 83.5 69.0 67.0 71.3 62.8
InternVL3-78B [77] 76.4 75.7 77.6 75.5 84.5 66.5 67.3 71.8 62.8
Qwen2.5-VL-7B [2] 74.3 77.4 69.7 72.1 88.8 55.5 68.8 83.5 54.0
Qwen2.5-VL-32B [2] 85.9 83.5 89.5 74.8 89.3 60.3 71.6 86.5 56.8
Qwen2.5-VL-72B [2] 84.8 90.8 80.9 79.4 88.8 70.0 76.3 84.3 68.3
Open-source Visual Grounded Reasoning Models
Pixel-Reasoner-7B [50] 80.6 83.5 76.3 72.9 86.0 60.3 66.9 80.0 54.3
DeepEyes-7B [76] 90.0 92.1 86.8 75.1 91.3 59.0 72.6 86.8 58.5
TreeVGR-7B 91.1 94.0 87.0 77.1 90.3 64.0 73.1 86.5 59.8
$\Delta$ v.s. Qwen2.5-VL-7B $\uparrow$ 16.8 $\uparrow$ 16.6 $\uparrow$ 17.3 $\uparrow$ 5.0 $\uparrow$ 1.5 $\uparrow$ 8.5 $\uparrow$ 4.3 $\uparrow$ 3.0 $\uparrow$ 5.8
Table 3: Comparison with state-of-the-art alternatives on V* Bench [66] and HRBench [63]. All results are self-collected. Best performances of visual grounded reasoning models are highlighted in bold.
Perception Reasoning
Overall OCR RS DT MO AD OCR DT MO AD
General Models
Qwen2.5-VL-7B [2] 42.3 87.6 32.7 83.0 27.3 30.0 72.0 62.0 28.7 23.0
Qwen2.5-VL-32B [2] 45.6 87.2 40.7 83.0 29.5 40.7 74.0 60.0 27.3 29.5
Qwen2.5-VL-72B [2] 43.7 90.8 34.0 87.0 27.9 30.6 74.0 61.0 26.7 25.5
LLaVA-OneVision-7B [23] 43.7 80.0 40.0 56.0 31.7 39.4 65.0 33.0 38.0 32.0
LLaVA-OneVision-72B [23] 48.7 79.2 50.7 67.0 37.9 40.0 76.0 41.0 38.7 39.3
InternVL3-8B [77] 47.9 83.6 49.3 75.0 34.5 36.9 70.0 44.0 40.0 37.0
InternVL3-38B [77] 51.0 85.6 56.0 71.0 42.6 40.0 77.0 45.0 47.3 35.0
InternVL3-78B [77] 52.3 87.6 54.7 77.0 42.6 36.6 76.0 56.0 46.0 40.3
Visual Grounded Reasoning Models
Pixel-Reasoner-7B [50] 49.7 89.6 52.0 86.0 38.9 30.9 71.0 72.0 46.0 32.5
DeepEyes-7B [76] 53.2 90.0 52.7 89.0 43.3 33.4 76.0 69.0 44.0 35.0
TreeVGR-7B 54.9 87.6 50.7 83.0 47.0 43.4 74.0 66.0 51.3 39.0
$\Delta$ v.s. Qwen2.5-VL-7B $\uparrow$ 12.6 – 0.0 $\uparrow$ 18.0 – 0.0 $\uparrow$ 19.7 $\uparrow$ 13.4 $\uparrow$ 2.0 $\uparrow$ 4.0 $\uparrow$ 22.6 $\uparrow$ 16.0
Table 4: Comparison with state-of-the-art alternatives on MME-RealWorld-Lite [72]. All results are self-collected. The best performance is highlighted in bold.
Correlation with Other Multimodal Benchmarks. We systematically compare our TreeBench with three existing multimodal benchmarks: AI2D [17], MathVista [37], and MMStar [7], in Figure 13, Figure 13, and Figure 13, respectively, to investigate potential performance correlations. Our analysis reveals a decoupling of performance characteristics. For instance, while GPT-4o-1120 [41] ranks among the top performers on TreeBench, it lags significantly behind alternatives on other benchmarks. This dissociation underscores the unique emphasis on “thinking with images” of our TreeBench.
Capability Benchmark Qwen2.5-VL-7B [2] TreeVGR-7B Qwen2.5-VL-72B [2]
Vision-centric question answering CV-Bench-2D [53] 74.1 76.9 $\uparrow$ 2.8 77.7
CV-Bench-3D [53] 72.6 77.6 $\uparrow$ 5.0 87.0
MMVP [54] 66.7 75.3 $\uparrow$ 8.6 66.7
[…]
POPE [26] 86.7 87.2 $\uparrow$ 0.5 84.9
HallusionBench [14] 48.2 50.1 $\uparrow$ 1.9 55.2
[…]
Table 5: Comparison with state-of-the-art alternatives on other multimodal benchmarks. Results are obtained from [16], otherwise are self-collected.
5.2 Main Properties of TreeVGR
Baselines and Benchmarks. We mainly compare our TreeVGR with state-of-the-art private models, including GPT-4o-1120 [41] and o3-0416 [43] from OpenAI, and Gemini-2.5-Flash-0520 [9] and Gemini-2.5-Pro-0605 [10] from Google DeepMind. Additionally, 8 representative open-source general models are incorporated, including LLaVA-OneVision series [23], Qwen2.5-VL series [2], and InternVL3 series [77]. Furthermore, two very recent visual grounded reasoning models are also included, i.e., DeepEyes [76] and Pixel-Reasoner [50], as both of them follow a “grounding then answering” pipeline, with the capability of “thinking with images”. Evaluations are mainly conducted on TreeBench, V* Bench [66], HR-Bench [63], and MME-RealWorld-Lite [72].
Results on High-Resolution Benchmarks. As demonstrated in Table 2, our TreeVGR-7B achieves 45.2 on our TreeBench, even surpasses Qwen2.5-VL-72B [2], which is comparable with open-source state-of-the-art InternVL3-78B [77]. Moreover, compared with visual grounded reasoning models, our TreeVGR not only achieves a higher overall performance, but also obtains larger mIoU, indicating its effectiveness in precisely localizing target objects. Additionally, in Table 3, TreeVGR also achieves open-source state-of-the-art on V* Bench [66] with 94.3 accuracy, which is even comparable with o3 [43]. On HR-Bench [63] and MME-RealWorld-Lite [72] illustrated in Table 3 and Table 4, respectively, our TreeVGR brings significant improvements over our base model, Qwen2.5-VL-7B [2].
Results on Other Multimodal Benchmarks. In Table 5, we compare our TreeVGR with its base model Qwen2.5-VL-7B [2] on a variety of conventional multimodal benchmarks. Specifically, we select CV-Bench [53], MMVP [54], and RealWorldQA [68], to evaluate vision-centric question-answering capabilities. MMBench [30], POPE [26], and HallusionBench [14] are selected for evaluating general VQA capabilities, and AI2D [17] and ChartQA [38] for comprehension with document and chart. We observe significant improvements in most cases, especially for vision-centric benchmarks. Notably, TreeVGR-7B achieves 75.3 on MMVP [54], even surpasses Qwen2.5-VL-72B [2] by a significant margin.
Rewards TreeBench V* [66] MME-RW [72]
Cold-Start $R_{\text{acc}}+R_{\text{format}}$ $R_{\text{IoU}}^{\text{R}}$ $R_{\text{IoU}}^{\text{P}}$ Acc mIoU Acc Acc
[…]
Table 6: Ablations of each component of our TreeVGR. “MME-RW” stands for MME-RealWorld-Lite [72], and “Acc” represents the multiple-choice accuracy. This improvement mainly comes from the training set, as many training samples from V* [66] are included in RL. The model tends to enumerate candidate boxes to obtain larger IoU recall, and fails to produce final answers.
5.3 Ablation Studies
The core contribution of TreeVGR is the traceable training pipeline, where the dual IoU reward $R_{\text{IoU}}$ is incorporated in conventional RL training. Therefore, we aim to evaluate the effectiveness of including this traceable term. As demonstrated in Table 6, we ablate each component of our TreeVGR, including the cost-start initialization and reward functions.
[…]
Figure 14: Mean response length with different IoU rewards. The precision term is crucial for alleviating the repetition problem.
[…]
6 Conclusion
This paper introduces TreeBench, a benchmark designed to rigorously evaluate visual grounded reasoning (VGR) or “thinking with images” in large multimodal models, and TreeVGR, a two-stage training framework that enhances VGR methods through traceable evidence supervision.
TreeBench addresses critical gaps in existing benchmarks by focusing on three principles: focused visual perception (identifying subtle targets in cluttered scenes), traceable evidence (quantifiable reasoning chains via bounding box annotations), and vision-centric second-order reasoning. Constructed through expert-driven annotation and multi-stage quality control, TreeBench features 405 high-difficulty visual question-answer pairs with precise bounding boxes, emphasizing small objects in real-world scenarios. It reveals the limitations of state-of-the-art models, e.g., OpenAI-o3 [43] scores 54.8%, while setting a new standard for assessing nuanced visual grounding, multi-step reasoning transparency, and cross-modal interaction.
TreeVGR advances VGR training through reinforcement learning guided by dual IoU rewards, which explicitly supervise bounding box generation to ensure both precision and recall. This approach enables explainable reasoning pathways and achieves significant improvements across benchmarks. TreeVGR achieves superior performance in just 5 epochs compared to 32 epochs for [76].
Limitation and future works. The current implementation of TreeVGR is based on a 7B parameter model, which may limit scalability compared to larger architectures. TreeBench contains only 405 rigorously curated question-answer pairs. Expanding the benchmark with additional samples across broader domains would further challenge model capabilities. Scaling up would be future work.